“BLACKFRIDAY” company is planning an expansion of its existing production capacity. The firm hired you as a consultant for the expansion project. Since you are a savvy project manager, you first decided to estimate the firm’s cost of capital based on the available data. Data: Tax Rate: 40% Bond: Coupon rate 12%, Maturity Years 15, Present value $1150 Preferred Stock: Dividend rate 10%, Par Value $100, Present Value $111 Common Stock: Market price $50, D0=$4.20, Dividend growth 5%, Beta 1.2, Treasury Bond yield 7%, Market risk premium 6%. When the firm uses Bond-yield+Premium method, the risk premium is 4%. Capital structure of “BLACKFRIDAY” is as follows; Debt 30%, Common Equity 60%, Preferred Stock 10% What are the components of the WACC you calculated above to be blamed for higher WACC? What can be done to reduce the WACC further?
Answer #1
Bond
Face Value $                     1,000.00
Coupon rate 12%
Coupon Payment = PMT $                        120.00
Maturity 15 Yrs
Present Value 1150
Cost of debt before tax = Rate = YTM = 6.21%
Cost of debt after tax = 6.21% x (1-40%) 3.72%
Preferred Stock
Par value $                        100.00
Dividend  rate 10%
Dividend   $                          10.00
Present Value $                        111.00
Cost of Preferred stock = Dividend/ PV = 10/111 9.01%
Equity
Cost of Equity = D1/P0  + G =( 4.20 x (1+5%)/$50 ) + 5% 13.82%
Cost of Equity = Rf + Beta x MRP = 7% + 1.2 x 6% 14.200%
Cost of Equity = Bond Yield + Risk Premium = 6.21% + 4% 10.21%
Average cost of Equity = (13.82% + 14.20% + 10.21%)/3 12.74%
WACC = We xRe + Wp x Rp + Wd x Rd aft tax
WACC = 60% x 12.74% + 10% x 9.01% + 30% x 3.72% 9.66%
Components of the WACC calculated above to be blamed for higher WACC are:
Risk Free Rate - When the Fed raises interest rates, the risk-free rate immediately increases and vice versa.
Tax Rate - Higher corporate taxes increase WACC, while lower taxes reduce WACC.
Longer Maturity of Debt - The longer the time to maturity on a firm’s debt, the longer it will be riskier and take for the full impact of higher rates to be felt.
Market Risk Premium -By  increasing volatility in the stock market will raise the risk premium demanded by investors
The most effective ways to reduce the WACC are (1) lower the cost of equity or (2) change the capital structure to include more debt.

A B 1000 0.12 =B4 B5 15 1150 =RATE(B7,B6,-38) =B9*60% 2 3 Bond 4 Face Value 5 Coupon rate 6 Coupon Payment = PMT 7 Maturity 8